OpenStudy (kkutie7):

(a) A stone is dropped from rest off of a cliff in the mountains. How long does it take for it to fall 100 m? You may assume that there is no air resistance. (b) How long does it take to fall from 100 m below its starting point to 200 m below its starting point? (c) Explain in words why the answers to (a) and (b) are not the same. (d) Assume now that the stone drops a total distance of h. Determine an expression (in term of whatever variables are needed) for its velocity after traveling that distance. Verify that the expression you obtained has the correct units.

OpenStudy (anonymous):

a) use h=0.5 *g*(t^2) where h is the height covered and g is the acceleration due to gravity and t i sthe required time... b) for be same use the formaula as in a but here h and t will be different from a d) the answers are different as the motion is under constant acceleration due to gravity in case a initial speed was 0 but after in case be i has already come down 100m and has acquired some initial velocity so the answers differ...

OpenStudy (kkutie7):

I solve the answer for (a). its 4.5s I used this formula \[r=r _{0}+v_{0}t+\frac{1}{2}at^{2}\] \[r=100,a=9.8,v_{0}=0\] I'm not getting a correct answer for b though.

OpenStudy (mtalhahassan2):

hey

OpenStudy (anonymous):

for b use r=200 r0=0 and v0=0 whatever t you obtain subtarct from it the value of t obtained in a

OpenStudy (kkutie7):

ok that works thanks

OpenStudy (anonymous):

welcome

OpenStudy (mtalhahassan2):

Hey Kkutie why you block me?

OpenStudy (kkutie7):

i didn't =(

OpenStudy (kkutie7):

for the last one I have \[h=0+0*t+\frac{1}{2}gt^{2}\] the answer is supposed to be \[-\sqrt{2hg}\]

OpenStudy (kkutie7):

I'm not sure how they got the answer. should h have to equal \[\frac{t^{2}}{2g}\]

OpenStudy (anonymous):

for determining velocity use v^2=u^2 +2gh

OpenStudy (kkutie7):

ok that makes way more sense. thank you again

OpenStudy (anonymous):

welcome